# algorithmic

### v2

@@ 4.2.2 Importance of Researcher-Labeler Agreement
High-quality labels are important for the final policy in an RLHF process. Early work on fine-tuning language models using RLHF noticed a mismatch between the researcher’s goals and the (paid) labeler’s actual labels (researcher-labeler disagreement). Ziegler et al. (2020) note that researchers agreed with each other about $60\%$ of the time (on 4-way comparisons, where random choice would result in $25\%$ agreement), while agreeing with labelers only $38\%$ or $46\%$ of the time (depending on the task). Stiennon et al. (2020) attempt to reduce these disagreements by maintaining a hands-on relationship with the labelers and thereby ensuring high researcher-labeler agreement. Concretely, they provide on-boarding with detailed instructions, keep an open channel of communication between researchers and labelers, and give feedback to the labelers. They evaluate the researcher-labeler agreement and reach an agreement rate of $77\%\pm 2\%$.
Perfect labels are often impossible due to the inherently subjective nature of the task. Returning to the example given by Stiennon et al. (2020), different researchers agreed with each other in only $73\%\pm 4\%$ of the cases. Ouyang et al. (2022) also report the agreement rates on a different task (instruction fine-tuning instead of summarization) and find that labelers agree with each other in $72.6\pm 1.5\%$ of the time, after a screening procedure that, amongst others, selects labelers that agree with researcher labels. Preferences can additionally be inconsistent between feedback types, as demonstrated by the finding of Bansal et al. (2024), which shows that preferences inferred from ratings and rankings significantly disagree for both human and AI annotators.
@@
The scale of the labeled data set required for effective training and refinement varies widely, impacting the quality of the resulting models. Studies have shown a broad range in data set sizes, from tens of labels in smaller studies (Jain et al., 2015) to hundreds in more complex scenarios (Christiano et al., 2017). Larger-scale applications may require thousands (Guan et al., 2021) or even millions of labels (Abramson et al., 2022), each bringing its own challenges in ensuring label accuracy and consistency. This variability in data set size underscores the need for rigorous label quality control measures across different scales. In smaller data sets, each label carries more weight, making accuracy and precision critical. Conversely, in larger data sets, the challenge lies in maintaining consistency and mitigating systematic biases that might emerge from the sheer volume of data.
Similarly, the labeling setting varies in the surveyed works, from author-provided feedback (Kim et al., 2023), over small in-person studies (Katz et al., 2021), to larger remote studies (Kim et al., 2023). Each setting provides unique challenges to ensure high-quality labels.
@@ 5 Reward Model Training
Figure 4: Highlighting the components of the RLHF loop discussed in this section.
In this section, we delve deeper into the process of reward model learning, which we briefly touched on in Section 2.3. We will discuss various aspects associated with this topic, namely the various human feedback models, utility (i.e., reward) learning per se, different reward model inputs, and how to increase feedback efficiency.
@@ 5.1 Human Feedback Model
The basic premise underlying the majority of approaches in RLHF is that human feedback is directly related to the reward function to be learned. To this end, the human feedback must first be captured in a sound mathematical framework that establishes the connection to the reward function. On a high level, one can break down (almost) all feedback types in Section 3.2 to a choice scenario: The human chooses one specific feedback option (label) from an available (possibly infinite) pool of possible feedback options (choice sets) [^2] This point of view goes back to the work of Jeon et al. (2020).. Here, the query that is made specifies the explicit contents of the choice set, e.g., if the query is to compare two trajectories, then the choice set consists of all possible outcomes for these two trajectories.
@@
(c $\mathbb{P}\left(\mbox{$c$ is chosen}\,|\,\mathcal{C}\right)\,\text{,}$ (4)
where $\mathcal{C}$ is the set of possible choices and $c\in\mathcal{C}$ one explicit choice. For the RLHF scenario, where the agent asks queries $q_{i}$ and the human gives labels $l_{i}$ as feedback (see Section 2.4), the choice set is specified by a function of the query. Formally, $\mathcal{C}=m(q)$ for some mapping $m$ that maps a query $q$ to the set of all possible candidate labels extractable from $q$ for the specific feedback type. For example, if the query is to rank a finite number of trajectories, then the choice set is the set of all possible rankings that can occur for the trajectories involved.
@@ 5.1.1 Boltzmann Distribution
(c $\mathbb{P}\left(\mbox{$c$ is chosen}\,|\,\mathcal{C}\right)=\mathbb{P}\left(c=\operatornamewithlimits{argmax}_{c\in\mathcal{C}}\ u_{c}+\epsilon_{c}\right)\,\text{.}$ (6)
@@
This is known as the Boltzmann distribution that also appears in a perhaps slightly modified version in various different subfields of ML and statistics. When restricting to discrete (choice) sets for $m(q)$, this distribution is also known as the multinomial logit model (Train, 2009) or Gibbs measure (Georgii, 2011), and as the Bradley-Terry model (Bradley & Terry, 1952) when the choice sets consist of pairs. All of these also have a link to the Plackett-Luce model (Luce, 1959; Plackett, 1975), which is a probability distribution on the space of total orders or rankings (see Alvo & Yu (2014) for details).
This model is often used for various reasons. A particularly compelling reason is the closed analytic form, which in turn makes it possible to obtain a closed form for the gradient with respect to the utilities. Another reason is that this model satisfies Luce’s axiom of choice (Plackett, 1975), which requires the probability of choosing an option from a pool of choice options not being affected by the presence or absence of other options in the pool. In this way, coherent decision-making is ensured, which, however, might be challenged as humans are likely not making fully rational decisions (see Section 4.2.1). Jeon et al. (2020) show that the usage of the Boltzmann distribution is justified by the principle of maximum entropy. More precisely, they show that it is the maximum entropy distribution over choices for a so-called satisficing human decision maker, i.e., one who is making in expectation a choice with an optimal reward up to some slack $\epsilon>0$.
@@
Note that $u_{l}$ depends on the return $R,$ so that we also may use $u_{l}(R)$ to emphasize this dependency. For the common case of pairwise trajectory comparisons, where for two trajectories $\tau_{1},\tau_{2}$ we obtain for the possible labels $l\in\{\tau_{1}\succ\tau_{2},\tau_{1}\prec\tau_{2}\}$ the respective utility by using the projection onto the preferred trajectory as the grounding function $G.$ Accordingly, the utility of the label represents essentially the utility of the preferred trajectory of that label, i.e., $\tau_{1}$ or $\tau_{2}$ in this case. As another example consider the case of e-stops feedback (see Section 3.2.2). Here, the possible labels $l$ provided by the user are $\texttt{STOP}_{t}$ and CONT encoding the stopping at time $t$ or the continuation of a trajectory. For the grounding function, one can define
$G(l)=\begin{cases}\tau,&l=\texttt{CONT},\\ \tau_{0:t}\tau_{t}\dots\tau_{t},&l=\texttt{STOP}_{t},\end{cases}$
where $\tau_{0:t}$ is the trajectory of $\tau$ trimmed to the stopping time $t$ and $\tau_{t}$ is the action-state pair at time $t.$ Table 1 in Jeon et al. (2020) provides an overview of the different grounding functions that lead to a specific feedback type. It is worth noting that one can also easily find a grounding function for the feedback type of a (partial) order over trajectories as considered, for instance, by Myers et al. (2021). Moreover, one can generalize this modeling approach by using (partial) segments instead of trajectories.
@@ 5.1.2 Human-Specific Rationality Coefficient
The Boltzmann distribution in (8) can be extended by a rationality coefficient $\beta\in[0,\infty)$ that reflects the precision of the human labeler [^3] This can be achieved by multiplying the utilities $u_{l}$ with $\beta$.:
@@
The higher $\beta$, the more (10) resembles a pointmass distribution modeling a highly rational human labeler (decision-maker) that is always able to identify the option the with highest utility, while the lower $\beta$, the more (10) resembles a uniform distribution modeling a highly irrational human labeler (decision-maker) acting purely at random. Without this extension, the commonly considered Boltzmann distribution (or Bradley-Terry model in the common case of pairwise comparisons) in (8) assumes a rationality coefficient of $1$. Ghosal et al. (2023) show in their experiments that the estimation of this coefficient can indeed positively influence reward learning. For the estimation, however, a calibration reward function is needed, as the rationality coefficient is otherwise not identifiable (Bengs & Hüllermeier, 2020). Similar findings are shown by Daniels-Koch & Freedman (2022), who model the rationality coefficient as a query-dependent function that might differ for the human labelers (see Section 5.1.6).
@@ 5.1.3 Alternative Utility Notions
Knox et al. (2024) show that the Boltzmann model does not generally lead to an identifiable reward function using (9) by presenting three concrete scenarios for which identification is not possible. The root cause of the non-identifiability is the usage of a trajectory’s return as the utility in (9). They, therefore, suggest using a trajectory’s regret as an alternative, which provably leads to identifiable rewards.
@@ 5.1.4 Human Feedback Models Beyond Boltzmann
While the human feedback model based on the Boltzmann distribution is the most widely used model nowadays, other models have also been considered in the literature. In particular, for the probability in (4) other models such as the Thurstone model (Wilson et al., 2012; Kupcsik et al., 2018; Bıyık et al., 2020), the ridge-noise model (Schoenauer et al., 2014), the binary model (Sugiyama et al., 2012) or mixed forms thereof (Wirth et al., 2016) have been considered. Of these models, only the Thurstone model (Thurstone, 1927) has a similar interpretation as the Boltzmann distribution based on perturbed utilities, only differing in the distribution of the perturbance random variables.
@@
Another possibility, which is particularly popular in theoretical work on RLHF (see Section 7), is the use of other functions on the right-hand sides of Eq. 8 than the exponential function. The concept is primarily used for pairwise comparisons of trajectories. It essentially states that the probability of the result of a pairwise comparison between two trajectories is the difference of their utility values under a so-called link function. More specifically, let $q=\{\tau_{1},\tau_{2}\}$ be the query to compare the trajectories $\tau_{1}$ and $\tau_{2}$, then, assuming a link function $\Phi:\mathbb{R}\to[0,1]$, one models the probability in (5) for $l$ representing a preference for $\tau_{1}$ as
@@
Note that the second property implies $\Phi(x)=1/2$ so that trajectories with the same utility also have the same chance of being selected. Any cumulative distribution function of a symmetric continuous random variable fulfills these two conditions. The two most common link functions that both fulfill the conditions are the linear link function given by
@@
Both are cumulative distribution functions: The linear link function is the cumulative distribution function of a continuous uniform distribution on $[0,1]$. In contrast, the logistic link function is the cumulative distribution function of a logistic distribution with location parameter 0 and scale parameter 1. Moreover, both are intensively studied in theoretical approaches (see Section 7.1), and the latter leads to (8) (when restricted on pairwise comparisons) and is a special case of the softmax function.
@@
Abramson et al. (2022) evaluate the usage of a generative model for learning from human preferences. More specifically, instead of assuming some underlying utility as in the Bradley-Terry model, they attempt to train a model to generate the human feedback (inter-temporal preferences in this case, see Section 3.2.1) and directly interpret this feedback as reward. However, they found that this empirically does not perform as well as the inter-temporal Bradley-Terry model.
@@ 5.1.6 Diverse Preferences
Bakker et al. (2022) investigate the first option by proposing to learn multiple reward functions, which can then be aggregated in arbitrary manners and even be utilized to find consensus among people with different preferences. The second is more commonly used, however: Xue et al. (2023a) learn a single reward function from multiple humans who may give diverse and inconsistent feedback, aiming to stabilize learning in spite of these inconsistencies using regularization, a consistency constraint, and ensembling. Similarly, Chhan et al. (2024) try to estimate the correct preference for pairwise trajectories directly by combining the users’ expressed preference labels instead of assuming individual reward functions. As a middle-ground between the single reward function and multiple ones, Myers et al. (2021) propose to learn a multimodal reward function that captures multiple people’s preferences and use a mixture model of Plackett-Luce models to represent the feedback more accurately.
@@ 5.1.7 Relaxation of the Markov Assumption
Most works assume that the human feedback is given based on a latent Markovian reward model, i.e., the return of a trajectory $\tau$ decomposes into a sum of independent rewards over state-action pairs (see (1)). Early et al. (2022) relax this assumption by dropping the need for the Markov property, such that the instantaneous reward might depend on hidden states. Similarly, Kim et al. (2023) avoid the Markov assumption by utilizing a transformer as the preference model. A similar effect may be achieved by learning a state representation with a recurrent network in which the rewards are Markov, similar to the approach taken by Hafner et al. (2023), but we are not aware of any work exploring this. Abramson et al. (2022) work in a non-Markovian setting as well by utilizing memory-augmented networks for both the policy and the reward model.

### v3

@@ 4.2.2 Importance of Researcher-Labeler Agreement
High-quality labels are important for the final policy in an RLHF process. Early work on fine-tuning language models using RLHF noticed a mismatch between the researcher’s goals and the (paid) labeler’s actual labels (researcher-labeler disagreement). Ziegler et al. (2020) note that researchers agreed with each other about $60\%$ of the time (on 4-way comparisons, where random choice would result in $25\%$ agreement), while agreeing with labelers only $38\%$ or $46\%$ of the time (depending on the task). Stiennon et al. (2020) attempt to reduce these disagreements by maintaining a hands-on relationship with the labelers, thereby ensuring high researcher-labeler agreement. Concretely, they provide on-boarding with detailed instructions, keep an open channel of communication between researchers and labelers, and give feedback to the labelers. They evaluate the researcher-labeler agreement and reach an agreement rate of $77\%\pm 2\%$.
Perfect labels are often impossible due to the inherently subjective nature of the task. Returning to the example given by Stiennon et al. (2020), different researchers agreed with each other in only $73\%\pm 4\%$ of the cases. Ouyang et al. (2022) also report the agreement rates on a different task (instruction fine-tuning instead of summarization) and find that labelers agree with each other in $72.6\pm 1.5\%$ of the time, after a screening procedure that, amongst others, selects labelers that agree with researcher labels. Preferences can additionally be inconsistent between feedback types, as demonstrated by the findings of Bansal et al. (2024), which show that preferences inferred from ratings and rankings significantly disagree for both human and AI annotators.
@@
The scale of the labeled data set required for effective training and refinement varies widely, impacting the quality of the resulting models. In practice, the required amount of data depends on many factors including domain, task complexity, and level of pretraining. For example, while LLM fine-tuning generally often requires tens of thousands of samples due to the complexity of the task, at the same time it can leverage the extensive prior knowledge of the pretrained LLM to reduce sample complexity compared to other tasks of similar complexity.
Studies have shown a broad range in data set sizes, from tens of labels in smaller studies (Jain et al., 2015) to hundreds in more complex scenarios (Christiano et al., 2017). Larger-scale applications may require thousands (Guan et al., 2021; Ouyang et al., 2022) or even millions of labels (Abramson et al., 2022), each bringing its own challenges in ensuring label accuracy and consistency.
This variability in data set size underscores the need for rigorous label quality control measures across different scales. In smaller data sets, each label carries more weight, making accuracy and precision critical. Conversely, in larger data sets, the challenge lies in maintaining consistency and mitigating systematic biases that might emerge from the sheer volume of data. The labeling setting varies in the surveyed works, from author-provided feedback (Kim et al., 2023), over small in-person studies (Katz et al., 2021), to larger remote studies (Kim et al., 2023). Each setting provides unique challenges to ensure high-quality labels.
@@ 5 Reward Model Training
Figure 4: RLHF diagram highlighting components discussed in this section.
In this section, we delve deeper into the process of reward model learning, which we briefly touched on in Section˜2.3. We will discuss various aspects associated with this topic, namely the various human feedback models, utility (i.e., reward) learning per se, different reward model inputs, and how to increase feedback efficiency.
@@ 5.1 Human Feedback Model
The basic premise underlying the majority of approaches in RLHF is that human feedback is directly related to the reward function to be learned. To this end, the human feedback must first be captured in a sound mathematical framework that establishes the connection to the reward function. On a high level, one can break down almost all feedback types in Section˜3.2 to a choice scenario: The human chooses one specific feedback option (label) from an available pool of possible feedback options (choice sets), which may be infinite [^3] This point of view goes back to the work of Jeon et al. (2020).. The query determines the explicit contents of the choice set. For example, if the query is to compare two trajectories, then the choice set consists of all possible outcomes for these two trajectories.
@@
$\mathbb{P}\left(\mbox{$c$ is chosen}\,|\,\mathcal{C}\right)\,\text{,}$ (4)
where $\mathcal{C}$ is the set of possible choices and $c\in\mathcal{C}$ one explicit choice. For the RLHF scenario, where the agent asks queries $q_{i}$ and the human gives labels $l_{i}$ as feedback (see Section˜2.4), the choice set is specified by a function of the query. Formally, $\mathcal{C}=m(q)$ for some mapping $m$ that maps a query $q$ to the set of all possible candidate labels extractable from $q$ for the specific feedback type. For example, if the query is to rank a finite number of trajectories, then the choice set is the set of all possible rankings that can occur for the trajectories involved.
@@ 5.1.1 Boltzmann Distribution
$\mathbb{P}\left(\mbox{$c$ is chosen}\,|\,\mathcal{C}\right)=\mathbb{P}\left(c=\operatornamewithlimits{argmax}_{c\in\mathcal{C}}\ u_{c}+\epsilon_{c}\right)\,\text{.}$ (6)
@@
This is known as the Boltzmann distribution that also appears in a perhaps slightly modified version in various different subfields of machine learning (ML) and statistics. When restricting to discrete (choice) sets for $m(q)$, this distribution is also known as the multinomial logit model (Train, 2009) or Gibbs measure (Georgii, 2011), and as the Bradley-Terry model (Bradley & Terry, 1952) when the choice sets consist of pairs. All of these also have a link to the Plackett-Luce model (Luce, 1959; Plackett, 1975), which is a probability distribution on the space of total orders or rankings (see Alvo & Yu (2014) for details).
This model is often used for various reasons. A particularly compelling reason is the closed analytic form, which in turn makes it possible to obtain a closed form for the gradient with respect to the utilities. Another reason is that this model satisfies Luce’s axiom of choice (Plackett, 1975), which requires the probability of choosing an option from a pool of choice options not being affected by the presence or absence of other options in the pool. In this way, coherent decision-making is ensured, which, however, might be challenged as humans likely do not make fully rational decisions (see Section˜4.2.1). Jeon et al. (2020) show that the usage of the Boltzmann distribution is justified by the principle of maximum entropy. More precisely, they show that it is the maximum entropy distribution over choices for a so-called satisficing human decision maker, i.e., one who is making a choice with an optimal reward in expectation up to some slack $\epsilon>0$.
@@
Note that $u_{l}$ depends on the return $R$, so that we also may use $u_{l}(R)$ to emphasize this dependency. For the common case of pairwise trajectory comparisons, where for two trajectories $\tau_{1},\tau_{2}$ we obtain for the possible labels $l\in\{\tau_{1}\succ\tau_{2},\tau_{1}\prec\tau_{2}\}$ the respective utility by using the projection onto the preferred trajectory as the grounding function $G$. Accordingly, the utility of the label represents essentially the utility of the preferred trajectory of that label, i.e., $\tau_{1}$ or $\tau_{2}$ in this case. As another example consider the case of e-stops feedback (see Section˜3.2.2). Here, the possible labels $l$ provided by the user are $\texttt{STOP}_{t}$ and CONT encoding the stopping at time $t$ or the continuation of a trajectory. For the grounding function, one can define
$G(l)=\begin{cases}\tau,&l=\texttt{CONT}\text{,}\\ \tau_{0:t}\tau_{t}\dots\tau_{t},&l=\texttt{STOP}_{t}\text{,}\end{cases}$
where $\tau_{0:t}$ is the trajectory of $\tau$ trimmed to the stopping time $t$, and $\tau_{t}$ is the action-state pair at time $t$. Table 1 in Jeon et al. (2020) provides an overview of the different grounding functions that lead to a specific feedback type. It is worth noting that one can also easily find a grounding function for the feedback type of a (partial) order over trajectories as considered, for instance, by Myers et al. (2021). Moreover, one can generalize this modeling approach by using (partial) segments instead of trajectories.
@@ 5.1.2 Human-Specific Rationality Coefficient
The Boltzmann distribution in (8) can be extended by a rationality coefficient $\beta\in[0,\infty)$ that reflects the precision of the human labeler:
@@
The higher $\beta$, the more (10) resembles a pointmass distribution modeling a highly rational human labeler that is always able to identify the option with highest utility. The lower $\beta$, the more it resembles a uniform distribution modeling a highly irrational human labeler acting purely at random. Without this extension, the commonly considered Boltzmann distribution (or Bradley-Terry model in the common case of pairwise comparisons) in (8) assumes a rationality coefficient of $1$. Ghosal et al. (2023) show in their experiments that the estimation of this coefficient can indeed positively influence reward learning. However, estimation requires a calibration reward function, as the rationality coefficient is otherwise not identifiable (Bengs & Hüllermeier, 2020). Similar findings are shown by Daniels-Koch & Freedman (2022), who model the rationality coefficient as a query-dependent function that might differ for the human labelers (see Section˜5.1.6).
@@ 5.1.3 Alternative Utility Notions
Knox et al. (2024) show that the Boltzmann model does not generally lead to an identifiable reward function using (9) by presenting three concrete scenarios for which identification is not possible. The root cause of the non-identifiability is the usage of a trajectory’s return as the utility in (9). They therefore suggest using a trajectory’s regret as an alternative, which provably leads to identifiable rewards.
@@ 5.1.4 Human Feedback Models Beyond Boltzmann
While the human feedback model based on the Boltzmann distribution is the most widely used model nowadays, other models have also been considered in the literature. In particular, for the probability in (4) other models such as the Thurstone model [^4] The probit model, as used in Bıyık et al. (2020), corresponds to Thurstone’s Case V. (Wilson et al., 2012; Kupcsik et al., 2018; Bıyık et al., 2020), the ridge-noise model (Schoenauer et al., 2014), the binary model (Sugiyama et al., 2012) or mixed forms thereof (Wirth et al., 2016) have been considered. Of these models, only the Thurstone model (Thurstone, 1927) has a similar interpretation as the Boltzmann distribution based on perturbed utilities, only differing in the distribution of the perturbance random variables.
@@
Another possibility, which is particularly popular in theoretical work on RLHF (see Section˜7), is the use of other functions on the right-hand sides of Eq.˜8 than the exponential function. The concept is primarily used for pairwise comparisons of trajectories. It essentially states that the probability of the result of a pairwise comparison between two trajectories is the difference of their utility values under a so-called link function. More specifically, let $q=\{\tau_{1},\tau_{2}\}$ be the query to compare the trajectories $\tau_{1}$ and $\tau_{2}$, then, assuming a link function $\Phi:\mathbb{R}\to[0,1]$, one models the probability in (5) for $l$ representing a preference for $\tau_{1}$ as
@@
Note that the second property implies $\Phi(0)=1/2$ so that trajectories with the same utility also have the same chance of being selected. Any cumulative distribution function of a symmetric continuous random variable fulfills these two conditions. The two most common link functions that both fulfill the conditions are the linear link function given by
@@
Both are cumulative distribution functions: The linear link function is the cumulative distribution function of a continuous uniform distribution on $[0,1]$. In contrast, the logistic link function is the cumulative distribution function of a logistic distribution with location parameter 0 and scale parameter 1. Moreover, both are intensively studied in theoretical approaches (see Section˜7.1), and the latter leads to (8) (when restricted on pairwise comparisons) and is a special case of the softmax function.
@@
Abramson et al. (2022) evaluate the usage of a generative model for learning from human preferences. More specifically, instead of assuming some underlying utility as in the Bradley-Terry model, they attempt to train a model to generate the human feedback (inter-temporal preferences in this case, see Section˜3.2.1) and directly interpret this feedback as reward. However, they found that this empirically does not perform as well as the inter-temporal Bradley-Terry model.
@@ 5.1.6 Diverse Preferences
Bakker et al. (2022) investigate the first option by proposing to learn multiple reward functions, which can then be aggregated in arbitrary manners and even be used to find consensus among people with different preferences. The second is more commonly used, however: Xue et al. (2024) learn a single reward function from multiple humans who may give diverse and inconsistent feedback, aiming to stabilize learning in spite of these inconsistencies using regularization, a consistency constraint, and ensembling. Similarly, Chhan et al. (2024) try to estimate the correct preference for pairwise trajectories directly by combining the users’ expressed preference labels instead of assuming individual reward functions. As a middle-ground between the single reward function and multiple ones, Myers et al. (2021) propose to learn a multimodal reward function that captures multiple people’s preferences and use a mixture model of Plackett-Luce models to represent the feedback more accurately.
@@
When dealing with diverse preferences, aggregating them via the mean may not always be suitable. Instead, after learning the individual preferences, one needs to address the important question of preference aggregation, which is studied in social choice theory (Brandt et al., 2016). Thus, a recent rapidly-developing research trend has started to investigate the tools developed in social choice in AI alignment (Conitzer et al., 2024; Sorensen et al., 2024; Mishra, 2023) in general and in RLHF (Dai & Fleisig, 2024; Zhong et al., 2024a; Park et al., 2024; Chakraborty et al., 2024b; Swamy et al., 2024; Lambert et al., 2023) in particular.
@@ 5.1.7 Relaxation of the Markov Assumption
Most works assume that the human feedback is given based on a latent Markovian reward model, i.e., the return of a trajectory $\tau$ decomposes into a sum of independent rewards over state-action pairs (see (1)). Early et al. (2022) relax this assumption by dropping the need for the Markov property, such that the instantaneous reward might depend on hidden states. Similarly, Kim et al. (2023) avoid the Markov assumption by using a transformer as the preference model. A similar effect may be achieved by learning a state representation with a recurrent network in which the rewards are Markov, similar to the approach taken by Hafner et al. (2023), but we are not aware of any work exploring this. Abramson et al. (2022) work in a non-Markovian setting as well by using memory-augmented networks for both the policy and the reward model.